OpenStudy (anonymous):

Solve w^4 -12w^2-2=0 w=

OpenStudy (angela210793):

replace w^4 with x^2 and u'll get x^2-12x-2=0 D=b^2-4ac=144-4*1*(-2)=152>0 x1=(-b+sqrD)/2a=(12+sqr152)/2=6+sqr38-->w=x^2--->w1=(6+sqr38)^2 x^2=(-b-sqrD)/2a=(12-sqr152)/2=6-sqr38--->w2=(6-sqr38)^2 I guess u know how to solve them now